In a group of 30 pupils, 20 are girls and 10 are boys. From these 30 pupils, 5 are chosen at random. In how many ways may they be chosen? And find the probability that there are exactly 3 girls in the group of 5 chosen.
Answer by stanbon(75887)

In a group of 30 pupils, 20 are girls and 10 are boys. From these 30 pupils, 5 are chosen at random. In how many ways may they be chosen?
Ans: 30C5 = 142506
=========================
And find the probability that there are exactly 3 girls in the group of 5 chosen.
------------------------------
Ans: [20C3*10C2]/30C5 = 513000.36/142506 = 0.36
=================================================
Cheers,
Stan H.